### v1

@@ Abstract
In multi-agent reinforcement learning (MARL), agents repeatedly interact across time and revise their strategies as new data arrives, producing a sequence of strategy profiles. This paper studies sequences of strategies satisfying a pairwise constraint inspired by policy updating in reinforcement learning, where an agent who is best responding in period $t$ does not switch its strategy in the next period $t+1$. This constraint merely requires that optimizing agents do not switch strategies, but does not constrain the other non-optimizing agents in any way, and thus allows for exploration. Sequences with this property are called satisficing paths, and arise naturally in many MARL algorithms. A fundamental question about strategic dynamics is such: for a given game and initial strategy profile, is it always possible to construct a satisficing path that terminates at an equilibrium strategy? The resolution of this question has implications about the capabilities or limitations of a class of MARL algorithms. We answer this question in the affirmative for mixed extensions of finite normal-form games.
@@ 1 Introduction
Game theory is a mathematical framework for studying strategic interaction between multiple self-interested agents, called players. Game theoretic models are pervasive in machine learning, appearing in multi-agent systems (Zhang et al., 2021; Gronauer and Diepold, 2022), single-agent reinforcement learning with multiple objectives (Hayes et al., 2022), in adversarial model training (Goodfellow et al., 2014, 2020; Bose et al., 2020), and throughout online learning theory in the form of worst-case performance guarantees (Cesa-Bianchi and Lugosi, 2006).
@@
In an $n$-player game, each player $i=1,\cdots,n,$ selects a strategy $x^{i}\in\mathcal{X}^{i}$ and receives a reward $R^{i}(x^{1},\dots,x^{n})$, which depends on the collective strategy profile ${\mathbf{x}}=(x^{1},\dots,x^{n})=:(x^{i},{\mathbf{x}}^{-i})$. Player $i$’s optimization problem is to best respond to the strategy ${\mathbf{x}}^{-i}$ of its counterparts, choosing $x^{i}\in\mathcal{X}^{i}$ to maximize $R^{i}(x^{i},{\mathbf{x}}^{-i})$. A strategy profile ${\mathbf{x}}_{*}=(x_{*}^{i})_{i=1}^{n}$ is called a Nash equilibrium if all players are simultaneously best responding to one another:
$R^{i}(x_{*}^{i},{\mathbf{x}}_{*}^{-i})\geq R^{i}(y^{i},{\mathbf{x}}_{*}^{-i}),\quad\forall y^{i}\in\mathcal{X}^{i},\,\forall i=1,\cdots,n.$
The related tasks of computing, approximating, and learning Nash equilibrium have attracted enduring attention in the broader machine learning community (Singh et al., 2000; Jafari et al., 2001; Hu and Wellman, 2003; Daskalakis et al., 2010; Nowé et al., 2012; Bravo et al., 2018; Flokas et al., 2020; Hsieh et al., 2021; Lu, 2023).
@@
In the field of multi-agent reinforcement learning (MARL), players use learning algorithms to iteratively revise their strategies in response to the observed history of play, producing a sequence $\{\widehat{{\mathbf{x}}}_{t}\}_{t\geq 1}$ in the set of strategy profiles ${\mathbf{X}}:=\mathcal{X}^{1}\times\cdots\times\mathcal{X}^{n}$. In the multi-agent setting, each player’s learning problem is complicated by numerous inherent challenges. First, there is a non-stationarity issue, as a given individual’s reward function changes whenever other agents revise their strategies, which subsequently prompts the individual to revise its own strategy. Second, key information about the game or about the strategies of other players may only be partially observable to an individual, necessitating estimation of various quantities. Analyzing the convergence properties of MARL algorithms can therefore be difficult, and the development of theoretical tools for such analysis is an important aspect of multi-agent learning theory.
A number of MARL algorithms are designed to approximate dynamical systems $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ on the set of strategy profiles ${\mathbf{X}}$ in which the next strategy for player $i$ is selected as $x^{i}_{t+1}=f^{i}({\mathbf{x}}_{t})$, where ${\mathbf{x}}_{t}=(x^{1}_{t},\dots,x^{n}_{t})$ is the strategy profile in period $t$. A representative sample of MARL algorithms of this type is offered in the next section. This approach facilitates analysis of the convergence behavior of the MARL algorithm, as it allows the analyst to separately consider the convergence of the idealized dynamical process $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ induced by the update functions $\{f^{i}\}_{i=1}^{n}$ on one hand and then consider the approximation of $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ by the MARL algorithm’s iterates $\{\widehat{{\mathbf{x}}}_{t}\}_{t\geq 1}$ on the other.
@@
Our primary interest in this work centers on update functions satisfying a quasi-rationality condition called satisficing: if $x^{i}$ is a best response to ${\mathbf{x}}^{-i}$, then $f^{i}(x^{i},{\mathbf{x}}^{-i})=x^{i}.$ That is, when an agent is already best responding, the update rule instructs the agent to continue using this strategy. This quasi-rationality constraint generalizes the well-studied best response update and is desirable for stability of the resulting dynamics, as it guarantees that Nash equilibrium strategy profiles are invariant under the dynamics. Moreover, the satisficing condition is only quasi-rational, in that it imposes no constraint on strategy updates when an agent is not best responding, and so allows for exploratory strategy updates. Such update rules are common in the literature on multi-agent learning theory (Blume, 1993; Marden and Shamma, 2012; Chasparis et al., 2013).
@@
The basic motivation of this paper is to better understand the capabilities of MARL algorithms that operate using the satisficing principle in selecting successive strategies, potentially augmented with random exploration when an agent is not best responding. Rather than studying a particular collection of strategy update functions, we study the problem on the level of sequences in ${\mathbf{X}}$, which allows us to implicitly account for experimental strategy updates. A sequence $({\mathbf{x}}_{t})_{t\geq 1}$ of strategy profiles is called a satisficing path if, for each player $i$ and time $t$, one has that $x^{i}_{t+1}=x^{i}_{t}$ whenever $x^{i}_{t}$ is a best response to ${\mathbf{x}}^{-i}_{t}$. The central research question of this paper is such:
@@
The resolution of this question has implications for the possible effectiveness of a class of MARL algorithms designed to seek Nash equilibrium. Indeed, the question has been answered in the affirmative for two-player normal-form games by Foster and Young (2006) and for $n$-player symmetric Markov games by Yongacoglu et al. (2023), and in both classes of games this has directly lead to MARL algorithms with convergence guarantees for approximating equilibria. A positive resolution of this question would remove a theoretical obstacle and establish that uncoordinated, distributed random search can effectively assist Nash-seeking algorithms achieve last-iterate convergence guarantees in a more general class of games than previously possible.
Our Contributions. In this work, we give a positive answer to the preceding research question for (the mixed extension of) any $n$-player normal-form game with finite action sets. That is, we show that for a finite $n$-player game $\Gamma$ and any initial strategy profile ${\mathbf{x}}_{1}$, there exists a satisficing path of finite length beginning at ${\mathbf{x}}_{1}$ and ending at a Nash equilibrium of the game $\Gamma$. This partially answers an open question posed by Yongacoglu et al. (2023).
@@
We prove our main result by analytically constructing a satisficing path from an arbitrary initial strategy profile to a Nash equilibrium. Our approach is somewhat counterintuitive, in that it does not attempt to seek Nash equilibrium by improving the performance of unsatisfied players (players who are not best responding at a given strategy profile), but rather by updating strategies in a way that increases the number of unsatisfied players at each round. This tactic heavily leverages the freedom afforded to unsatisfied players to explore their strategy space and avoids the well-observed challenges of cyclical strategy revision that occurs when agents attempt to best respond to their counterparts.
@@
Notation. We let $\mathbb{P}$ and $\mathbb{E}$ denote probability and expectation, respectively. For a finite set $A$, we let $\mathbb{R}^{A}$ denote the vector space $\mathbb{R}^{|A|}$ with components indexed by elements of $A$. We let $\Delta_{A}$ denote the set of probability measures over a set $A$. For $n\in\mathbb{N}$, we let $[n]:=\{1,2,\dots,n\}$. For a point $x$, the Dirac measure centered at $x$ is denoted $\delta_{x}$. Agent indices are typically superscripts, while time/iteration indices are typically subscripts. Boldface characters are reserved for multi-agent quantities. When discussing a fixed agent $i$, the remaining collection of agents are called $i$’s counterparts or counterplayers.
Related Work
Beginning with fictitious play (Brown, 1951), a vast number of MARL algorithms have been proposed for iterative strategy adjustment while playing a game under various assumptions on observability of counterplayer strategies. The most widely studied class of algorithms of this type involve each player running a no-regret algorithm on its own stream of rewards. Fictitious play and its descendants, such as stochastic fictitious play (Hofbauer and Sandholm, 2002) and generalized weakened fictitious play (Leslie and Collins, 2006) are special cases of this class and have been extensively studied. Although the convergence behavior of fictitious play and its variants has been studied in several game models, including normal-form games and Markov games, convergence results are typically available only for games exhibiting special structure, such as zero-sum rewards or other advantageous properties amenable to analysis (Hofbauer and Sandholm, 2002; Baudin and Laraki, 2022; Sayin et al., 2022a, b).
The formalization of the fictitious play property is especially relevant to this paper. A game is said to have the fictitious play property if the empirical frequencies of strategies converge to a Nash equilibrium from any initialization. Examples of games with and without the fictitious play property date to (Robinson, 1951) and (Shapley, 1964), respectively, and the identification of games with and without this property was viewed as an important research question (Monderer and Sela, 1996; Monderer and Shapley, 1996a).
@@
The convergence properties of the fictitious play algorithm are intimately connected to those of best response dynamics, a full information dynamical system evolving in continuous time where the evolution rule for player $i$’s strategy is governed by its best response multi-function. By harnessing such connections, convergence results for fictitious play and a number of other MARL algorithms have been obtained by analyzing the dynamical systems induced by specific update rules (Benaïm et al., 2005; Leslie and Collins, 2005; Swenson et al., 2018b).
@@
A second research direction relevant to the present work involves (im)possibility results for strategic dynamics defined by strategy update functions, taking the form $x^{i}_{t+1}=f^{i}({\mathbf{x}}_{t})$ in discrete time or an analogous form in continuous time. The question of when such dynamics converge to Nash equilibrium has received persistent attention. In the case of deterministic strategy updates, Hart and Mas-Colell (2003) studied strategic dynamics in continuous time and showed that if the strategy update functions, analogous to $f^{i}$ above, satisfy regularity conditions as well as a desirable property called uncoupledness, by which $f^{i}$ cannot depend on the reward functions of $i$’s counterplayers, then the resulting dynamics are not Nash convergent in general. These results were recently generalized by Milionis et al. (2023), who obtained impossibility results in both continuous and discrete time while requiring only continuity of the deterministic dynamical system. Additional possibility and impossibility results were presented by Babichenko (2012), who studied strategic dynamics in a different setting, where players observe only their own actions and not the actions of their counterplayers.
Passing from deterministic strategic dynamics to stochastic strategic dynamics, a number of positive results were obtained by incorporating exogenous randomness into one’s strategy update, along with finite recall of recent play (Hart and Mas-Colell, 2006; Foster and Young, 2006; Germano and Lugosi, 2007).
In the regret testing algorithm of Foster and Young (2006), players revise their strategies according to whether or not their most recent strategy met a satisfaction criterion: if $x^{i}_{t}$ performed within $\epsilon$ of the optimal performance against ${\mathbf{x}}^{-i}_{t}$, player $i$ continues using it and picks $x^{i}_{t+1}=x^{i}_{t}$. Otherwise, player $i$ experiments and selects $x^{i}_{t+1}$ according to a uniformly positive probability distribution over $\mathcal{X}^{i}$. Conditional strategy updates similar to this have appeared in several other works, such as (Chien and Sinclair, 2011; Candogan et al., 2013; Chasparis et al., 2013), and the regret testing algorithm has been extended in several ways (Germano and Lugosi, 2007; Arslan and Yüksel, 2017).
A game is said to have the satisficing paths property if it admits a finite length satisficing path ending at equilibrium with an arbitrary initial strategy profile. As we discuss in the next section, satisficing paths can be interpreted as a natural generalization of best response paths. Consequently, the problem of identifying games that have the satisficing paths property is a theoretically relevant question analogous to characterizing exact potential games (Monderer and Shapley, 1996b) or determining when a game has the fictitious play property.
@@
The concept of satisficing paths was first formalized by Yongacoglu et al. (2023) in the context of multi-state Markov games, where it was shown that $n$-player symmetric Markov games have the satisficing paths property and this fact could be used to produce a convergent MARL algorithm. However, the core idea of satisficing paths appeared earlier, before this formalization: in the convergence analysis of the regret testing algorithm in (Foster and Young, 2006), it was shown that two-player normal-form games have the satisficing paths property, though this terminology was not used.
2 Model
A finite, $n$-player normal-form game $\Gamma$ is described by a list
$\Gamma=(n,{\mathbf{A}},{\mathbf{r}}),$
@@ 2 Model
where $n$ is the number of players, ${\mathbf{A}}=\mathbb{A}^{1}\times\cdots\times\mathbb{A}^{n}$ is a finite set of action profiles, and ${\mathbf{r}}=(r^{i})_{i\in[n]}$ is a collection of reward functions, where $r^{i}:{\mathbf{A}}\to\mathbb{R}$ describes the reward of player $i$ as a function of the action profile. The $i^{\rm th}$ component of ${\mathbf{A}}$ is player $i$’s action set $\mathbb{A}^{i}$.
@@
Description of play. Each player $i\in[n]$ selects its action $a^{i}$ according to a probability vector $x^{i}\in\Delta_{\mathbb{A}^{i}}$. That is, $a^{i}\sim x^{i}$. The vector $x^{i}$ is called player $i$’s mixed strategy, and we denote player $i$’s set of mixed strategies by $\mathcal{X}^{i}:=\Delta_{\mathbb{A}^{i}}$. Players are assumed to select their actions simultaneously, or at least without observing one another’s actions, and the collection of actions $\{a^{i}:i\in[n]\}$ is assumed to be mutually independent. The set of mixed strategy profiles is denoted ${\mathbf{X}}:=\mathcal{X}^{1}\times\cdots\mathcal{X}^{n}$ and corresponds to the set of product measures on ${\mathbf{A}}$. After the action profile is selected, each player $i$ receives reward $r^{i}(a^{1},\dots,a^{n})$.
@@
Player $i$’s performance criterion is its expected reward, which depends jointly on its mixed strategy and the strategies of $i$’s counterplayers. For each strategy profile ${\mathbf{x}}\in{\mathbf{X}}$, player $i$’s expected reward is
@@
Definition 2.2.
For $\epsilon\geq 0$, a strategy profile ${\mathbf{x}}_{*}=(x^{i}_{*},{\mathbf{x}}^{-i}_{*})\in{\mathbf{X}}$ is called an $\epsilon$-Nash equilibrium if, for every $i\in[n]$,
$R^{i}(x^{i}_{*},{\mathbf{x}}^{-i}_{*})\geq R^{i}(x^{i},{\mathbf{x}}^{-i}_{*})-\epsilon\quad\forall x^{i}\in\mathcal{X}^{i}.$
Putting $\epsilon=0$ in the definitions above, one recovers the classical definitions of best responding and Nash equilibrium. For any $\epsilon\geq 0$, the set of $\epsilon$-best responses to a specified strategy ${\mathbf{x}}^{-i}$ is denoted ${\rm BR}^{i}_{\epsilon}({\mathbf{x}}^{-i})\subseteq\mathcal{X}^{i}$.

### v2

@@ Abstract
In multi-agent reinforcement learning (MARL) and game theory, agents repeatedly interact and revise their strategies as new data arrives, producing a sequence of strategy profiles. This paper studies sequences of strategies satisfying a pairwise constraint inspired by policy updating in reinforcement learning, where an agent who is best responding in one period does not switch its strategy in the next period. This constraint merely requires that optimizing agents do not switch strategies, but does not constrain the non-optimizing agents in any way, and thus allows for exploration. Sequences with this property are called satisficing paths, and arise naturally in many MARL algorithms. A fundamental question about strategic dynamics is such: for a given game and initial strategy profile, is it always possible to construct a satisficing path that terminates at an equilibrium? The resolution of this question has implications about the capabilities or limitations of a class of MARL algorithms. We answer this question in the affirmative for normal-form games. Our analysis reveals a counterintuitive insight that reward deteriorating strategic updates are key to driving play to equilibrium along a satisficing path.
@@ 1 Introduction
Game theory is a mathematical framework for studying strategic interaction between self-interested agents, called players. In an $n$-player normal-form game, each player $i=1,\cdots,n,$ selects a strategy $x^{i}\in\mathcal{X}^{i}$ and receives a reward $R^{i}(x^{1},\dots,x^{n})$, which depends on the collective strategy profile ${\mathbf{x}}=(x^{1},\dots,x^{n})=:(x^{i},{\mathbf{x}}^{-i})$. Player $i$’s optimization problem is to best respond to the strategy ${\mathbf{x}}^{-i}$ of its counterparts, choosing $x^{i}\in\mathcal{X}^{i}$ to maximize $R^{i}(x^{i},{\mathbf{x}}^{-i})$. Game theoretic models are pervasive in machine learning, appearing in fields such as multi-agent systems [21], multi-objective reinforcement learning [24], and adversarial model training [7], among many others.
@@
In multi-agent reinforcement learning (MARL), players use learning algorithms to revise their strategies in response to the observed history of play, producing a sequence $\{\widehat{{\mathbf{x}}}_{t}\}_{t\geq 1}$ in the set of strategy profiles ${\mathbf{X}}:=\mathcal{X}^{1}\times\cdots\times\mathcal{X}^{n}$. Due to the coupled reward structure of multi-agent systems, each player’s learning problem involves a moving target: since an individual’s reward function depends on the strategies of the others, strategy revision by one agent prompts other agents to revise their own strategies. Convergence analysis of MARL algorithms can therefore be difficult, and the development of tools for such analysis is an important aspect of multi-agent learning theory.
A strategy profile $(x_{*}^{i})_{i=1}^{n}$ is called a Nash equilibrium if all players simultaneously best respond to one another. Nash equilibrium is a concept of central importance in game theory, and the tasks of computing, approximating, and learning Nash equilibrium have attracted enduring attention in theoretical machine learning [47, 27, 14, 42, 17, 26, 31]. Convergence to equilibrium strategies has long been a predominant, but not unique, design goal in MARL [53]. In this paper, we study mathematical structure of normal-form games with the twin objectives of (i) better understanding the capabilities or limitations of existing MARL algorithms and (ii) producing insights for the design of new MARL algorithms.
@@
A number of MARL algorithms approximate dynamical systems $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ on the set of strategy profiles ${\mathbf{X}}$ in which the next strategy for player $i$ is selected as $x^{i}_{t+1}=f^{i}({\mathbf{x}}_{t})$, where ${\mathbf{x}}_{t}=(x^{1}_{t},\dots,x^{n}_{t})$ is the strategy profile in period $t$. A sampling of such algorithms will be offered shortly. This approach facilitates analysis of the algorithm, as one separately considers the convergence of $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ induced by the update functions $\{f^{i}\}_{i=1}^{n}$, on one hand, and the approximation of $\{{\mathbf{x}}_{t}\}_{t\geq 1}$ by the algorithm’s iterates $\{\widehat{{\mathbf{x}}}_{t}\}_{t\geq 1}$ on the other. In this work, we consider update functions that satisfy a quasi-rationality condition called satisficing: when an agent is best responding, the update rule instructs the agent to continue using this strategy. That is, if $x^{i}$ is a best response to ${\mathbf{x}}^{-i}$, then $f^{i}(x^{i},{\mathbf{x}}^{-i})=x^{i}.$ This quasi-rationality constraint generalizes the best response update and is desirable for stability of the resulting dynamics, as it guarantees that Nash equilibria are invariant under the dynamics. Moreover, the satisficing condition is only quasi-rational, in that it imposes no constraint on strategy updates when an agent is not best responding, and so allows for exploratory strategy updates. Update rules that incorporate exploratory random search when a strategy is deemed unsatisfactory are common in MARL theory [6, 32, 11, 34].
@@
Our goal is to better understand the capabilities/limitations of MARL algorithms that use the satisficing principle to select successive strategies, potentially augmented with random exploration when an agent is not best responding. Examples include [19, 20, 33, 12, 10, 1] and [52]. Instead of studying a particular collection of strategy update functions, we abstract the problem to the level of sequences in ${\mathbf{X}}$, which allows us to implicitly account for experimental strategy updates. A sequence $({\mathbf{x}}_{t})_{t\geq 1}$ of strategy profiles is called a satisficing path if, for each player $i$ and time $t$, one has that $x^{i}_{t+1}=x^{i}_{t}$ whenever $x^{i}_{t}$ is a best response to ${\mathbf{x}}^{-i}_{t}$. The central research question of this paper is such:
@@
Since many MARL algorithms operate using the satisficing principle (or otherwise approximate processes that involve satisficing update rules, e.g. [48]), the resolution of this question has implications for the effectiveness of such MARL algorithms. Indeed, the question has been answered in the affirmative for two-player normal-form games by [19] and for $n$-player symmetric Markov games by [52], and in both classes of games this has directly lead to MARL algorithms with convergence guarantees for approximating equilibria. In addition to removing a theoretical obstacle, positive resolution of this question would establish that uncoordinated, distributed random search can effectively assist Nash-seeking algorithms to achieve last-iterate convergence guarantees in a more general class of games than previously possible.
@@
Contributions. We give a positive answer to the question above: for any finite $n$-player game $\Gamma$ and any initial strategy profile ${\mathbf{x}}_{1}$, there exists a satisficing path beginning at ${\mathbf{x}}_{1}$ and ending at a Nash equilibrium of $\Gamma$. This partially answers an open question posed by [52]. We prove this result by analytically constructing a satisficing path from an arbitrary initial strategy profile to a Nash equilibrium. Our approach is somewhat counterintuitive, in that it does not attempt to seek Nash equilibrium by improving the performance of unsatisfied players (players who are not best responding at a given strategy profile), but by updating strategies in a way that increases the number of unsatisfied players at each round. This tactic leverages the freedom afforded to unsatisfied players to explore their strategy space and avoids the challenge of cyclical strategy revision that occurs when agents attempt to best respond to their counterparts [37]. This insight provides a new approach to MARL algorithm design beyond the well-structured settings considered in prior work.
@@
Notation. We let $\Delta_{A}$ denote the set of probability measures over a set $A$. For $n\in\mathbb{N}$, we let $[n]:=\{1,2,\dots,n\}$. For a point $x$, the Dirac measure centered at $x$ is denoted $\delta_{x}$. When discussing a fixed agent $i$, the remaining collection of agents are called $i$’s counterparts or counterplayers.
Related Work.
@@
A vast number of MARL algorithms have been proposed for iterative strategy adjustment while playing a game. The most widely studied class of algorithms of this type involve each player running a no-regret algorithm on its own stream of rewards. The celebrated fictitious play algorithm [9] and its descendants are special cases of this class. Although the convergence behavior of fictitious play and its variants has been studied extensively, convergence results are typically available only for games exhibiting special structural properties amenable to analysis [25, 29, 4, 45, 46]. Indeed, the convergence properties of fictitious play are intimately connected to those of best response dynamics, a full information dynamical system evolving in continuous time where the evolution rule for player $i$’s strategy is governed by its best response multi-function. By harnessing such connections, convergence results for fictitious play and a number of other MARL algorithms have been obtained by analyzing the dynamical systems induced by specific update rules [5, 28, 49].
@@
A related line of research considers strategic dynamics defined by strategy update functions, taking the form $x^{i}_{t+1}=f^{i}({\mathbf{x}}_{t})$ in discrete time or an analogous form in continuous time. In the case of deterministic strategy updates, [22] studied strategic dynamics in continuous time and showed that if the strategy update functions, analogous to $f^{i}$ above, satisfy regularity conditions as well as a desirable property called uncoupledness, by which $f^{i}$ cannot depend on the reward functions of $i$’s counterplayers, then the resulting dynamics are not Nash convergent in general. These results were recently generalized by [38]. Additional possibility and impossibility results were presented by [2], who studied strategic dynamics in a different setting, where players do not observe counterplayer strategies. Under stochastic strategic dynamics, a number of positive results were obtained by incorporating exogenous randomness into one’s strategy update, along with finite recall of recent play [23, 19, 20]. In the regret testing algorithm of [19], players revise their strategies according to whether or not their most recent strategy met a satisfaction criterion: if $x^{i}_{t}$ performed within $\epsilon$ of the optimal performance against ${\mathbf{x}}^{-i}_{t}$, player $i$ continues using it and picks $x^{i}_{t+1}=x^{i}_{t}$. Otherwise, player $i$ experiments and selects $x^{i}_{t+1}$ according to a probability distribution over $\mathcal{X}^{i}$. Conditional strategy updates similar to this have appeared in several other works, such as [12, 10, 11], and the regret testing algorithm has been extended in several ways [20, 1].
@@
A game is said to have the satisficing paths property if every initial strategy profile is connected to some equilibrium by a satisficing path. As we discuss in the next section, satisficing paths can be interpreted as a natural generalization of best response paths. Consequently, the problem of identifying games that have the satisficing paths property is a theoretically relevant question analogous to characterizing potential games [41] or determining when a game has the fictitious play property [39, 40]. The concept of satisficing paths was first formalized in [52] in the context of multi-state Markov games, where it was shown that $n$-player symmetric Markov games have the satisficing paths property and this fact could be used to produce a convergent MARL algorithm. However, the core idea of satisficing paths appeared earlier, before this formalization: in the convergence analysis of the regret testing algorithm in [19], it was shown that two-player normal-form games have the satisficing paths property, though this terminology was not used. These earlier works made no claims about the existence of paths in general-sum $n$-player games, which is the focus of this paper.
2 Normal-form games
@@
A finite, $n$-player normal-form game $\Gamma$ is described by a list $\Gamma=(n,{\mathbf{A}},{\mathbf{r}}),$ where $n$ is the number of players, ${\mathbf{A}}=\mathbb{A}^{1}\times\cdots\times\mathbb{A}^{n}$ is a finite set of action profiles, and ${\mathbf{r}}=(r^{i})_{i\in[n]}$ is a collection of reward functions, where $r^{i}:{\mathbf{A}}\to\mathbb{R}$ describes the reward of player $i$ as a function of the action profile. The $i^{\rm th}$ component of ${\mathbf{A}}$ is player $i$’s action set $\mathbb{A}^{i}$.
@@
Description of play. Each player $i\in[n]$ selects a probability vector $x^{i}\in\Delta_{\mathbb{A}^{i}}$ and then selects its action $a^{i}$ according to $a^{i}\sim x^{i}$. The vector $x^{i}$ is called player $i$’s mixed strategy, and we denote player $i$’s set of mixed strategies by $\mathcal{X}^{i}:=\Delta_{\mathbb{A}^{i}}$. Players are assumed to select their actions without observing one another’s actions, and the collection of actions $\{a^{i}:i\in[n]\}$ is assumed to be mutually independent. The set of mixed strategy profiles is denoted ${\mathbf{X}}:=\mathcal{X}^{1}\times\cdots\mathcal{X}^{n}$. After the action profile ${\mathbf{a}}=(a^{1},\dots,a^{n})$ is selected, each player $i$ receives reward $r^{i}({\mathbf{a}})$.
@@
Player $i$’s performance criterion is its expected reward, defined for each strategy profile ${\mathbf{x}}\in{\mathbf{X}}$ as
@@
Definition 2.
For $\epsilon\geq 0$, a strategy profile ${\mathbf{x}}_{*}=(x^{i}_{*},{\mathbf{x}}^{i}_{*})\in{\mathbf{X}}$ is called an $\epsilon$-Nash equilibrium if, for every player $i\in[n]$, $x^{i}_{*}$ is an $\epsilon$-best response to ${\mathbf{x}}^{-i}_{*}$.
Putting $\epsilon=0$ above, one recovers the classical definitions of best responding and Nash equilibrium. For any $\epsilon\geq 0$, the set of $\epsilon$-best responses to a strategy ${\mathbf{x}}^{-i}$ is denoted ${\rm BR}^{i}_{\epsilon}({\mathbf{x}}^{-i})\subseteq\mathcal{X}^{i}$.